# Point Neuron Learning: A New Physics-Informed Neural Network Architecture

### v1

@@ 4.4 Circular microphone array placement
Figure 4 depicts the performance metrics NMSE (1l) and MAC (1m) of the proposed method, harmonics-based method, and PINN method as a function of frequency. Overall, the proposed method outperforms the other two methods for all frequencies, especially when the frequency is higher than 1100 Hz. Specifically, based on the NMSE result as shown in Fig. 4a, (i) the proposed method has the lowest NMSE for all frequencies, and it slightly increases from $-22$ dB in $100$ Hz to $-13$ dB in $2000$ Hz, (ii) the PINN method can roughly estimate the sound field under $1100$ Hz, while above $1100$ Hz the NMSE reaches 00 dB, and (iii) the NMSE for the harmonics-based method fluctuates between $-10$ dB to $10$ dB, and peaks are caused by the zeros of the spherical Bessel function. For small target regions, the harmonics-based method is reliable, but for larger target regions, more basis functions are used to estimate the sound field, and more Bessel zeros are involved in the estimation, resulting in poor performance of the harmonic-based method.
@@ 5 Conclusion
Appendix A Derivation of equation (1g)
@@
$\mathcal{P}^{v}_{q}=w_{v}\frac{D_{v}}{D_{q}^{v}}e^{ik(D_{q}^{v}-D_{v})}.$ (1o)
For updating point neuron weights, from (1ea), we have
$\frac{\partial\mathcal{L}}{\partial w_{v}^{*}}=\frac{\partial\big{(}\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big{)}}{\partial w_{v}^{*}}+\lambda\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}.$ (1p)
@@
$+$ $\frac{\partial\big{(}({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big{)}}{\partial{\mathcal{P}_{q}}^{*}}\frac{\partial{\mathcal{P}_{q}}^{*}}{\partial w_{v}^{*}}\Big{)},$ (1qa)
@@
$\frac{\partial\big{(}({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big{)}}{\partial{\mathcal{P}_{q}}}=$ $({\mathcal{P}_{q}-P_{q}})^{*},$ (1ra)
$\frac{\partial\mathcal{P}_{q}}{\partial w_{v}^{*}}=$ $0,$ (1rb)
$\frac{\partial\big{(}({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big{)}}{\partial{\mathcal{P}_{q}}^{*}}=$ $({\mathcal{P}_{q}-P_{q}}),$ (1rc)
$\frac{\partial{\mathcal{P}_{q}}^{*}}{\partial w_{v}^{*}(n)}=$ $\sum^{V}_{v^{\prime}=1}\frac{\partial{\mathcal{P}^{v^{\prime}}_{q}}^{*}}{\partial w_{v}^{*}}=\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}.$ (1rd)
Substituting (1r) into (1q)
$\frac{\partial\big{(}\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big{)}}{\partial w_{v}^{*}}=\sum^{Q}_{q=1}({\mathcal{P}_{q}-P_{q}})\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}.$ (1s)
@@
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}=$ $\sum^{V}_{v^{\prime}=1}\Big{(}\frac{\partial(|w_{v^{\prime}}|)}{\partial w_{v}^{*}}\Big{)}=\frac{1}{2}\Big{(}\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}+i\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}\Big{)},$ (1ta)
@@
$\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}=$ $\frac{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}{\sqrt{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}^{2}+\operatorname{\mathbb{I}m}\{w_{v}^{*}\}^{2}}}=\mathrm{cos}\theta_{v},$ (1ua)
$\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}=$ $\frac{\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}{\sqrt{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}^{2}+\operatorname{\mathbb{I}m}\{w_{v}^{*}\}^{2}}}=\mathrm{sin}\theta_{v}.$ (1ub)
Substituting (1u) into (1t)
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}=\frac{1}{2}e^{i\theta_{v}}.$ (1v)
Based on (1p), (1s), and (1v), we have
$\frac{\partial\mathcal{L}}{\partial w_{v}^{*}}=\sum^{Q}_{q=1}({\mathcal{P}_{q}-P_{q}})\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}\,+\frac{1}{2}\lambda e^{i\theta_{v}}.$ (1w)
Appendix B Derivation of equation (1i)
@@
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=\frac{\partial(\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}})}{\partial B_{v}^{x}}+\lambda\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial B_{v}^{x}}.$ (1x)
@@
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial B_{v}^{x}}=0,$ (1y)
@@
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=$ $\frac{\partial\big{(}\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big{)}}{\partial B_{v}^{x}}$ (1za)
@@
$+$ $\frac{\partial\big{(}({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big{)}}{\partial\mathcal{P}_{q}^{*}}\frac{\partial\mathcal{P}_{q}^{*}}{\partial D_{v}}\frac{\partial D_{v}}{\partial B_{v}^{x}}\Big{)}$ (1zb)
@@
$\frac{\partial\big{(}({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big{)}}{\partial\mathcal{P}_{q}}=$ $({\mathcal{P}_{q}-P_{q}})^{*},$ (1aaa)
@@
$=$ $(\frac{\partial\mathcal{P}_{q}}{\partial D_{q}^{v}})^{*},$ (1aac)
@@
$=$ $(\frac{\partial\mathcal{P}_{q}^{*}}{\partial D_{v}})^{*},$ (1aae)
$\frac{\partial D_{v}}{\partial B_{v}^{x}}=$ $\frac{B_{v}^{x}}{D_{v}},$ (1aaf)
$\frac{\partial D_{q}^{v}}{\partial B_{v}^{x}}=$ $\frac{(B_{v}^{x}-x_{q})}{D_{q}^{v}}.$ (1aag)
Substituting (1aa) into (1z), we have
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=$ $\sum_{q=1}^{Q}~{}2\operatorname{\mathbb{R}e}\Bigg{\{}({\mathcal{P}_{q}-P_{q}})^{*}w_{v}e^{ik(D_{v}^{q}-D_{v})}\frac{D_{v}}{D_{v}^{q}}$ (1aba)
$\times$ $\Big{(}\frac{-(ikD_{v}-1)}{{D_{v}}^{2}}B_{v}^{x}+\frac{(ikD_{v}^{q}-1)}{{D_{v}^{q}}^{2}}(B_{v}^{x}-x_{q})\Big{)}\Bigg{\}}.$ (1abb)
The derivations of (1j) and (1k) are similar to the derivation of (1i).

### v2

@@ 4.4 Circular microphone array placement
Figure 4 depicts the performance metrics NMSE (12) and MAC (13) of the proposed method, harmonics-based method, and PINN method as a function of frequency. Overall, the proposed method outperforms the other two methods for all frequencies, especially when the frequency is higher than 1100 Hz. Specifically, based on the NMSE result as shown in Fig. 4a, (i) the proposed method has the lowest NMSE for all frequencies, and it slightly increases from 22-22 dB in $100$ Hz to 13-13 dB in $2000$ Hz, (ii) the PINN method can roughly estimate the sound field under $1100$ Hz, while above $1100$ Hz the NMSE reaches 0 dB, and (iii) the NMSE for the harmonics-based method fluctuates between 10-10 dB to $10$ dB, and peaks are caused by the zeros of the spherical Bessel function. For small target regions, the harmonics-based method is reliable, but for larger target regions, more basis functions are used to estimate the sound field, and more Bessel zeros are involved in the estimation, resulting in poor performance of the harmonic-based method.
@@ 5 Conclusion
Appendix A Derivation of equation (7)
@@
$\mathcal{P}^{v}_{q}=w_{v}\frac{D_{v}}{D_{q}^{v}}e^{ik(D_{q}^{v}-D_{v})}.$ (15)
For updating point neuron weights, from (5a), we have
$\frac{\partial\mathcal{L}}{\partial w_{v}^{*}}=\frac{\partial\big(\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big)}{\partial w_{v}^{*}}+\lambda\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}.$ (16)
@@
$+$ $\frac{\partial\big(({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big)}{\partial{\mathcal{P}_{q}}^{*}}\frac{\partial{\mathcal{P}_{q}}^{*}}{\partial w_{v}^{*}}\Big),$ (17a)
@@
$\frac{\partial\big(({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big)}{\partial{\mathcal{P}_{q}}}=$ $({\mathcal{P}_{q}-P_{q}})^{*},$ (18a)
$\frac{\partial\mathcal{P}_{q}}{\partial w_{v}^{*}}=$ $0,$ (18b)
$\frac{\partial\big(({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big)}{\partial{\mathcal{P}_{q}}^{*}}=$ $({\mathcal{P}_{q}-P_{q}}),$ (18c)
$\frac{\partial{\mathcal{P}_{q}}^{*}}{\partial w_{v}^{*}(n)}=$ $\sum^{V}_{v^{\prime}=1}\frac{\partial{\mathcal{P}^{v^{\prime}}_{q}}^{*}}{\partial w_{v}^{*}}=\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}.$ (18d)
Substituting (18) into (17)
$\frac{\partial\big(\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big)}{\partial w_{v}^{*}}=\sum^{Q}_{q=1}({\mathcal{P}_{q}-P_{q}})\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}.$ (19)
@@
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}=$ $\sum^{V}_{v^{\prime}=1}\Big(\frac{\partial(|w_{v^{\prime}}|)}{\partial w_{v}^{*}}\Big)=\frac{1}{2}\Big(\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}+i\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}\Big),$ (20a)
@@
$\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}=$ $\frac{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}}{\sqrt{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}^{2}+\operatorname{\mathbb{I}m}\{w_{v}^{*}\}^{2}}}=\mathrm{cos}\theta_{v},$ (21a)
$\frac{\partial(|w_{v}|)}{\partial\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}=$ $\frac{\operatorname{\mathbb{I}m}\{w_{v}^{*}\}}{\sqrt{\operatorname{\mathbb{R}e}\{w_{v}^{*}\}^{2}+\operatorname{\mathbb{I}m}\{w_{v}^{*}\}^{2}}}=\mathrm{sin}\theta_{v}.$ (21b)
Substituting (21) into (20)
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial w_{v}^{*}}=\frac{1}{2}e^{i\theta_{v}}.$ (22)
Based on (16), (19), and (22), we have
$\frac{\partial\mathcal{L}}{\partial w_{v}^{*}}=\sum^{Q}_{q=1}({\mathcal{P}_{q}-P_{q}})\frac{D_{v}}{D_{q}^{v}}e^{-ik(D_{q}^{v}-D_{v})}\,+\frac{1}{2}\lambda e^{i\theta_{v}}.$ (23)
Appendix B Derivation of equation (9)
@@
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=\frac{\partial(\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}})}{\partial B_{v}^{x}}+\lambda\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial B_{v}^{x}}.$ (24)
@@
$\frac{\partial(\rVert\boldsymbol{w}\rVert_{1})}{\partial B_{v}^{x}}=0,$ (25)
@@
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=$ $\frac{\partial\big(\sum^{Q}_{q=1}{|\mathcal{P}(\boldsymbol{x}_{q})-P(\boldsymbol{x}_{q})|^{2}}\big)}{\partial B_{v}^{x}}$ (26a)
@@
$+$ $\frac{\partial\big(({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big)}{\partial\mathcal{P}_{q}^{*}}\frac{\partial\mathcal{P}_{q}^{*}}{\partial D_{v}}\frac{\partial D_{v}}{\partial B_{v}^{x}}\Big)$ (26b)
@@
$\frac{\partial\big(({\mathcal{P}_{q}-P_{q}})({\mathcal{P}_{q}-P_{q}})^{*}\big)}{\partial\mathcal{P}_{q}}=$ $({\mathcal{P}_{q}-P_{q}})^{*},$ (27a)
@@
$=$ $(\frac{\partial\mathcal{P}_{q}}{\partial D_{q}^{v}})^{*},$ (27c)
@@
$=$ $(\frac{\partial\mathcal{P}_{q}^{*}}{\partial D_{v}})^{*},$ (27e)
$\frac{\partial D_{v}}{\partial B_{v}^{x}}=$ $\frac{B_{v}^{x}}{D_{v}},$ (27f)
$\frac{\partial D_{q}^{v}}{\partial B_{v}^{x}}=$ $\frac{(B_{v}^{x}-x_{q})}{D_{q}^{v}}.$ (27g)
Substituting (27) into (26), we have
$\frac{\partial\mathcal{L}}{\partial B_{v}^{x}}=$ $\sum_{q=1}^{Q}~2\operatorname{\mathbb{R}e}\Bigg\{({\mathcal{P}_{q}-P_{q}})^{*}w_{v}e^{ik(D_{v}^{q}-D_{v})}\frac{D_{v}}{D_{v}^{q}}$ (28a)
$\times$ $\Big(\frac{-(ikD_{v}-1)}{{D_{v}}^{2}}B_{v}^{x}+\frac{(ikD_{v}^{q}-1)}{{D_{v}^{q}}^{2}}(B_{v}^{x}-x_{q})\Big)\Bigg\}.$ (28b)
The derivations of (10) and (11) are similar to the derivation of (9).